Ringmeup Inc. had net income of $160,000 for the year ended December 31, 2019. At the beginning of the year, 39,000 shares of common stock were outstanding. On May 1, an additional 18,000 shares were issued. On December 1, the company purchased 4,600 shares of its own common stock and held them as treasury stock until the end of the year. No other changes in common shares outstanding occurred during the year. During the year, Ringmeup paid the annual dividend on the 6,000 shares of 4.25%, $100 par value preferred stock that were outstanding the entire year.

Required:
Calculate basic earnings per share of common stock for the year ended December 31, 2019. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Earnings per share = ?
Answer #1

Weighted average share = (39000*4/12+57000*7/12+52400*1/12) = 50616.66 shares

Preferred dividend = 6000*100*4.25% = 25500

Earning per share = (Net income-Preferred dividend)/Weighted average share = (160000-25500)/50616.66 = 2.66 per share